The net cash flows of Advantage Leasing for the next 3 years are $42,000, $49,000 and $64,000 respectively, after which the growth rate will be a constant 2% with a WACC of 8%. What is the present value of the terminal value?


Select one:
a. 51,822
b. 55,967
c. 59,259
d. 60,444
e. None of these
Answer #1

Cash flow at the end of third year = 64,000


Terminal value at the end of third year is given by C*(1+g)/(r-g) = 64,000*(1+2%)/(8% - 2%) = 1088000. This terminal value is the value at the end of third year. Its present value is given by 1088000/(1+r)^3 = 1088000/(1+8%)^3 = 863689.5.


Hence, the answer is e. None of these